5. Dueto adverse economic circumstances and poor management, Sultan Company has negotiated a restructuring of its P10,000,000 note payable to Kudarat Bank. Kudarat Bank has agreed to reduce the face value of the note to P8,000,000 and extend the due date three years from the date of restructuring. However the interest rate was increased from 15% to 2190, The restructuring will occur on December 31, 2016. There is no unpaid interest on the restructured loan at this time. The tax rate-is 35% How much is the gain on extinguishment of debt for the year 2016? (Round off present value factors to four decimal places) P2,000,000 P1,999,848 P904,224 0
Answer #1
PV (Carrying Value)of Note payable extinguished 10000000
PV of the new Note payable (PV of after-tax interest payments PLUS PV of Return of Face value at end of Year 3
ie. ((8000000*21%)*2.3353)+(8000000*0.6812) 9372904
PV Gain on extinguishment of debt 627096
NOTE: After-tax cost of new debt=21%(1-35%)=13.65%       &
P/A,13.65%,3 yrs.=2.3353
P/F, 13.65%, Yr.3= 0.6812